Calculate the magnitude of the electric field at the center of a square 42.5cm on a side if one corner is occupied by a −38.8μC charge and the other three are occupied by −27.4μCcharges.
Answer #1

electric field due to two corner charges having charge - 27.4 uC are equal and opposite hence they cancel each other. so the net electric field is only due to - 27.4 and - 38.8 uC

E = k ( q1 - q2) / ( a^2/2)

E = 9*10^9* 10^-6* ( 38.8 - 27.4) / (0.425^2 / 2)

E = 1.136*10^6 N/C

=====

Comment in case any doubt, will reply for sure.. Goodluck